A researcher wants to estimate the numbers of hours that 5 years old children spend watching television. A sample of 50 five-year old children was observed to have a mean viewing time of 5 hours. The population is normally distributed with a population standard deviation of 0.5 hours. Find the 95% confidence interval of the population mean.


1
Expert's answer
2022-05-12T08:59:05-0400

The critical value for "\alpha = 0.05" is "z_c = z_{1-\alpha/2} = 1.96."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-z_c\dfrac{\sigma}{\sqrt{n}}, \bar{x}+z_c\dfrac{\sigma}{\sqrt{n}})"

"=(5-1.96\dfrac{0.5}{\sqrt{50}}, 5+1.96\dfrac{0.5}{\sqrt{50}})"

"=(4.861, 5.139)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "4.861 < \mu < 5.139," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(4.861, 5.139)."